OpenStudy (anonymous):

What is the volume of a cone with a radius of 6 cm and a height of 12 cm? Use 3.14 for pi. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

V=3.14*R^2*H/3 V=3.1486^2*12/3=452.39cm^3

OpenStudy (anonymous):

Okay, what about; A cylindrical well has a radius of 8 feet and a height of 13 feet. What is the volume of the well?

OpenStudy (anonymous):

In other words a cylinder?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

V=3.14R^2H V=3.14*8^2*13=2613.81ft^3

OpenStudy (anonymous):

Thanks. One more?

OpenStudy (anonymous):

Ok...

OpenStudy (anonymous):

What is the volume of a sphere with a diameter of 8 meters?

OpenStudy (anonymous):

No, thats why I'm asking. @Isaiah.Feynman

OpenStudy (anonymous):

V=4/3*3.14R^3 V=4/3*3.14*4^3=268.08

OpenStudy (anonymous):

Thanks so much! And yes, I would.

OpenStudy (anonymous):

Ur welcome